Two forces equal to 2p and p respectively act on a partical. If the first force is doubled and the sec9nd be increased by 12N the direction if the resultant un unaltered . F8nd the value of p

Grade:12th pass

1 Answers

Vikas TU
7 years ago

From the given parameters,

 F1net = root(P+ 4p2 + 4p2) = underroot(9p2)

F2net = underroot(4p+ (2p + 20)2 + 4p.(2p+20)) = underroot(16p2 + 160p +400)

Both the forces would be equal as given,

Hence, F1net = F2net

7p2 +160p +400 = 0

p = - 40/28 = -10/7 N

and also

p = - 100 N

These would be the forces acting in oppsote direction.